t would cost 78.85 to buy three tickets in japan plus two tickets in switzerland. three tickets in switzerland plus two tickets in japan would cost 74.80. how much does an average movie ticket cost in each of these countries
Answer by mananth(16949)

Japan x
Switzerland y

3 x + 2 y = 78.85 .............1
2 x + 3 y = 74.8 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
-9 x -6 y = -236.55
4 x 6 y = 149.6
Add the two equations
-5 x = -86.95
/ -5
x = 17
plug value of x in (1)
3 x + 2 y = 78.85
52.17 + 2 y = 78.85
2 y = 78.85 -52.17
2 y = 26.68
y = 13
Japan 17.39/ticket
Switzerland 13.34/ticket
m.ananth@hotmail.ca